### v2

# Simple Controller Design to Achieve Iso-Damping Robustness: Non-Iterative Data-Driven Approach Based on Fractional-Order Reference Model

Abstract
This study proposes a simple controller design approach to achieve a class of robustness, the so-called iso-damping property. The proposed approach can be executed using only one-shot input/output data. An accurate mathematical model of a controlled plant is not required. The model-reference control problem is defined to achieve the desired closed-loop specifications, including the iso-damping, and the reference model is designed on the basis of fractional-order calculus. The optimization problem for the model-reference control is formulated using the one-shot input/output data while considering the bounded-input bounded-output (BIBO) stability from a bounded reference input to a bounded output. The iso-damping robust controller is obtained by solving the optimization problem. The representative advantages of the proposed approach over the conventional methods are the simplicity, practicality, and reliability from the viewpoint of the unnecessity of the plant model and explicit consideration of the BIBO stability from a bounded reference input to a bounded output. Numerical examples demonstrate the validity of the proposed approach.
Index Terms:
Fractional-order control; Data-driven control; Iso-damping; Parameter tuning; Fictitious reference signal; Optimization.
I Introduction
I-A Motivation
Robustness is a fundamental requirement for control systems in practice, since the controlled plant has various uncertainties and characteristic changes. Numerous studies have been conducted to achieve robust control [1]. A conventional approach to ensure robustness is 1) modeling the nominal plant and assuming the range of the uncertainty and 2) designing the controller based on the model-based control theory to work well in the pre-specified uncertainty range. However, this approach fails if the plant model including uncertainty is not appropriate [2]; modeling the nominal plant and the uncertainty relies on the expert’s knowledge, increasing the development cost. Therefore, establishing a practical, simple, and reliable framework for robust controller design is an important challenge in the control engineering community.
Among the various types of uncertainties, gain variation is often found in practical control systems (e.g., load changes in robotic systems [3]). Gain variation adversely affects the control performance and even causes instability. Thus, the robustness to plant gain variation is an indispensable requirement for practical control systems. Recently, the iso-damping property [4] has garnered a great deal of attention. The iso-damping property is a special class of robustness: the invariance of the overshoot of the setpoint response under plant gain variation. Hence, the iso-damping property is highly desirable for practical control systems, enhancing the reliability and safety.
Remark 1.
If the Bode phase plot of the open-loop transfer function is locally flat around the gain crossover frequency, such a special characteristic is called the flat-phase property [4]. A feedback control system with the flat-phase property achieves the iso-damping robustness [4]. $\blacksquare$
I-B Related works
Control systems involving a fractional-order calculus (FC) has generated a great deal of attention: fractional-order (FO) system theory (e.g., stabilization of FO systems [5, 6, 7, 8, 9], stability tests [10, 11, 12, 13], the generalized KYP lemma for singular FO systems [14], the FC-based gradient algorithm [15]). FC is a mathematical framework generalizing traditional integer-order (IO) calculus, which considers non-integer order differentiations and integrations. Focusing on its rich expression ability and generality, Magin et al. have discussed the benefit of exploiting the idea of FC in the cybernetics field [16]. In particular, FO control is a powerful and effective approach to achieving good control performance and robustness [17]. For example, the Bode’s ideal transfer function (BITF) is an FO transfer function with the ideal flat-phase specification; it is often utilized for controller design to achieve the iso-damping robustness [18]. FC-based robust control is one important direction of the exploitation of FC in the cybernetics field.
Various approaches have been proposed to achieve iso-damping on the basis of FO control. These methods aim to tune some design parameters of the fixed structure controller, such as the FO proportional-integral-derivative (PID) controller, so as to realize the flat-phase property. The analytical conditions for the flat-phase have been derived for a special class of plants and controllers: the FO-PID controller for simplified FO plants with time delay [19], the FO-PID controller for stable first-order plus time delay (FOPTD) plant [20], and the FO-PID controller with FO lowpass filter [21]. The graphical approach (e.g., plotting the feasible parameter region or achievable performance specification) is an effective way to determine controller parameters satisfying the analytical flat-phase condition. As for the graphical approach, one study has utilized the Bode phase plot [22], one has proposed a flat-phase curve [23], and a novel procedure to tune the FO-[PID] controller has been established for the FOPTD plant [24]. It should be noted that both the analytical and graphical approaches require accurate mathematical models and parameters of the actual controlled plant. However, the necessity of the plant model and parameters imposes heavy burdens on designers and causes the technique to be fail under modeling errors. Moreover, in most cases, the analytical and graphical policies are derived for very simple plants such as the FOPTD system. Thus, not only the plant model but also its bold approximation is necessary to employ such a method. Controller design based on the approximated model often causes severe performance degradation due to the gap between the actual plant and the reduced one for controller design.
Numerical optimization-based approaches have been examined for a wide variety of plants in the context of parameter tuning of FO control systems (e.g., the differential evolution algorithm [25], the wolf based elephant herding method [26]). As for realizing the flat-phase specification, a novel metaheuristics algorithm has been developed for FO-PID controller tuning [27]. Nevertheless, the conventional optimization-based design requires an accurate plant model and parameters for evaluating the objective function.
Recently, controller design techniques based on the data rather than the mathematical model of the plant have attracted a great deal of attention: direct data-driven control (D3C) [28, 29, 30]. Exploiting the D3C approach to design an FO control system can overcome the drawbacks of the abovementioned challenges, since these problems are mainly raised by the needs for accurate plant models and parameters. Several studies have employed frequency response data for designing FO controllers achieving the iso-damping property [31, 32]. However, obtaining the frequency response is cumbersome and often requires a special experimental setup.
Model-reference (MR) D3C is a simple and practical framework for achieving the desired closed-loop characteristics while avoiding burdensome frequency response computation. Various MR-D3C techniques have been examined for IO control systems: iterative feedback tuning (IFT) [33], virtual reference feedback tuning (VRFT) [34, 35], correlation-based tuning [36], fictitious reference iterative tuning (FRIT) [37], to name a few. Many studies have reported the effectiveness of MR-D3C in the IO control framework [38, 39, 40, 41]. For example, the VRFT approach has been employed for the vehicle yaw rate control [42, 43].
On the other hand, to the best of our knowledge, few studies have explored an MR-D3C approach for FO control. The IFT technique has been employed to tune the FO-PID controller for a hydraulic actuator [44]. However, the tuning scheme presented in [44] requires multiple experiments, imposing heavy burdens on designers. The FO-PID controller has been tuned on the basis of the VRFT [45, 46, 47]. Nevertheless, the conventional VRFT does not consider closed-loop stability. Although one study has tuned the FO-PID controller based on the MR-D3C approach, it does not consider the iso-damping robustness [48]. Consequently , to the best of our knowledge, no studies have proposed simple and practical FO MR-D3C techniques especially for achieving the iso-damping property.
I-C Contribution and novelty
This study presents a novel controller design scheme for a discrete-time single-input single-output (SISO) linear time-invariant (LTI) system. The proposed approach does not require a mathematical model of the controlled plant. The present approach can be executed using only one-shot input/output data collected by a closed-loop experiment. The resultant controller due to the proposed approach can achieve iso-damping: the overshoot amount of the closed-loop system is almost invariant under the plant gain variation. In the present approach, the robust controller design problem is defined as the MR-D3C problem using the reference model a priori chosen by the designer. The reference model is designed on the basis of an FO system, defining the desired characteristics of the closed-loop system such as the flat-phase property. The MR-D3C problem is reformulated as an optimization problem based on a fictitious reference signal (the fictitious reference signal is computed using the one-shot input/output data). The loss function to be minimized reflects the information of the closed-loop poles from a reference input to an output. The solution of the minimization problem yields a controller achieving the desired closed-loop characteristic such as the iso-damping specified by the reference model. Numerical examples demonstrate the validity of the proposed approach.
Compared with conventional approaches focusing on iso-damping robust control, the contribution and novelty can be summarized as follows:
  1. (1)
    (Simplicity) The proposed approach does not require the mathematical model of the actual controlled plant, providing freedom from costly and burdensome modeling procedures. Therefore, the present approach is simpler than model-based approaches such as analytical and graphical approaches [19, 20, 21, 22, 23, 24] and numerical optimization based on the plant model [18, 27, 25].
  2. (2)
    (Optimality) The existing analytical and graphical approaches are developed for very simple plants, requiring not only the plant model but also model approximations. The gap between the actual plant and the approximated model design causes the performance degradation. On the other hand, the proposed technique does not require the model reduction since the plant model itself is unnecessary. Therefore, the present approach is expected to provide superior (i.e., more optimal) controller than the approximated-model-based techniques.
  3. (3)
    (Reduced burdens on designers) The appropriate controller is automatically obtained via solving the optimization problem in the proposed approach. The manual trial-and-error requiring the expert’s skill is unnecessary. Moreover, the proposed tuning scheme is executed on the basis of only one-shot input/output data, unlike the traditional FO MR-D3C technique requiring multiple datasets [44]. Thus, the proposed approach significantly reduces the burdens on the designer.
  4. (4)
    (Reliability) Unlike the conventional FO MR-D3C techniques [45, 46, 47], information regarding the closed-loop pole is explicitly reflected in the proposed optimization problem for the controller design. This feature allows the present technique to explicitly consider the bounded-input bounded-output (BIBO) stability of the closed-loop system from a bounded reference input to a bounded output. Therefore, the proposed design approach yields a reliable controller.
  5. (5)
    (Generality) Most analytical approaches are developed for a special class of plants and controllers, as discussed above. On the other hand, the proposed approach is applicable to a wide variety of discrete-time SISO LTI plants and controllers.
Owing to contributions (1)–(5), we can easily incorporate useful and important robustness, i.e., the iso-damping, into various control systems. This study contributes to improve the safety, reliability, and performance of a wide variety of control systems in practice. Moreover, this study has the importance from the viewpoint of the incorporation of FC into the cybernetics field. Specifically, this study is a pioneering work of the exploration the interdisciplinary area between data-driven control and the FO system theory.
I-D Structure of this paper
This paper is organized as follows. Section II summarizes some preliminaries and states the problem addressed in this study. Section III describes the proposed controller design approach. We show the numerical studies to demonstrate the validity of the proposed approach in Section IV. The discussion is given in Section V. Section VI concludes the paper.
II Preliminaries
The symbols $\mathbb{R}$ and $\mathbb{R}_{+}$ denote the sets of real numbers and strictly positive real numbers, respectively. The set of the real-valued matrices with the dimension of $m\times n$ is represented as $\mathbb{R}^{m\times n}$ (we write $\mathbb{R}^{n\times 1}$ as $\mathbb{R}^{n}$). For a vector $v\in\mathbb{R}^{n}$, $\mathrm{Tpl}\left(\cdot\right)\colon\mathbb{R}^{n}\mapsto\mathbb{R}^{n\times n}$ is defined as
$\mathrm{Tpl}(\it{v})\coloneqq\begin{bmatrix}v_{1}&0&0&\ddots&0\\ \vdots&v_{1}&0&\ddots&0\\ v_{n-2}&\ddots&v_{1}&0&\vdots\\ v_{n-1}&v_{n-2}&\ddots&\ddots&0\\ v_{n}&v_{n-1}&v_{n-2}&\ldots&v_{1}\\ \end{bmatrix}.$ (1)
Here, $v_{j}\;(j=1,2,\ldots n)$ is the $j$-th component of $v$. For an $n$-length sequence $v=\{v_{k}\}_{k=0}^{n-1}\;(1\leq n\leq\infty)$, $\mathrm{vec}(v)\coloneqq\begin{bmatrix}v_{0}&v_{1}&\ldots&v_{n-1}\end{bmatrix}^{\top}$ and $\mathrm{Tpl}\circ\mathrm{vec}\coloneqq\mathrm{Tpl}\left(\mathrm{vec}(\cdot)\right)$. For a matrix $M\in\mathbb{R}^{n\times n}$, $\Lambda\left(M\right)$ denotes the set of all eigenvalues of $M$ including their algebraic multiplicity. The spectral radius of $M$ is represented as $\rho\left(M\right)$, i.e., $\rho(M)\coloneqq\max_{\lambda\in\Lambda(M)}\left\lvert\lambda\right\rvert$.
[…]
Let $G\left(z\right)$ be a proper rational discrete-time IO SISO LTI system and let $\{g_{k}\}_{k=0}^{\infty}$ be the impulse response sequence of $G\left(z\right)$. Then, the output $y=\{y_{k}\}_{k=0}^{\infty}$ of $G\left(z\right)$ to the input $u=\{u_{k}\}_{k=0}^{\infty}$ is $y_{t}=g_{t}\ast u_{t}\;(t=0,1,2,\ldots)$. Here, the symbol $\ast$ denotes the convolution, i.e., $g_{t}\ast u_{t}\coloneqq\sum_{n=0}^{t}g_{\tau}u_{t-\tau}=\sum_{n=0}^{t}g_{t-\tau}u_{\tau}$ [49]. We sometimes denote the input/output relation as $y=G(z)u$.
The $p$-norm $\left\lVert v\right\rVert_{p}$ of $v\in\mathbb{R}^{n}$ is defined as $\left\lVert v\right\rVert_{p}\coloneqq\left(\sum_{k=1}^{n}\left\lvert v_{j}\right\rvert^{p}\right)^{\frac{1}{p}}$ for $1\leq p<\infty$ and $\left\lVert v\right\rVert_{\infty}\coloneqq\sup_{j}\left\lvert v_{j}\right\rvert$. Similarly, the $l_{p}$-norm $\left\lVert u\right\rVert_{p}$ of a real-valued signal $u=\{u_{k}\}_{k=0}^{\infty}$ is defined as $\left\lVert u\right\rVert_{p}\coloneqq\left(\sum_{k=1}^{\infty}\left\lvert u_{j}\right\rvert^{p}\right)^{\frac{1}{p}}$ for $1\leq p<\infty$ and $\left\lVert u\right\rVert_{\infty}\coloneqq\sup_{j}\left\lvert u_{j}\right\rvert$.
II-A The flat-phase property and iso-damping
Consider the feedback control system shown in Fig. 1, where $L\left(s\right)$ is a continuous-time SISO LTI transfer function and $K$ is the static gain. Let $\omega_{c}\in\mathbb{R}_{+}$ be such that $\left\lvert L\left(j\omega_{c}\right)\right\rvert=1$. The flat-phase property can be characterized as
$\left.\frac{d}{ds}\left(\angle L(s)\right)\right|_{s=j\omega_{c}}=0$ (2)
where $\angle L\left(s\right)$ represents the argument of $L\left(s\right)$ [18]. If $L\left(s\right)$ in Fig. 1 has the flat-phase property, the closed-loop system shown in Fig. 1 has a special robustness: the overshoot amount of the output $y$ to the setpoint reference input $r$ is invariant under the variation of $K$, namely, the iso-damping [4]. The iso-damping property is practically useful since the plant gain variation often occurs in many control systems.
Remark 2.
A typical choice of $L\left(s\right)$ realizing the flat-phase property is the BITF $L_{Bode}\left(s\right)$ defined as
$L_{Bode}\left(s\right)=\left(\frac{\omega_{c}}{s}\right)^{\gamma}$ (3)
where $\gamma\in\mathbb{R}$. Note that the BITF is an FO transfer function since the FO integrator/differentiator becomes $s^{\alpha}\;(\alpha\in\mathbb{R})$ in the Laplace domain ($\alpha<0$ for integrator, $\alpha>0$ for differentiator). This fact clearly shows that the FO control framework is effective to realize the iso-damping property, owing to its high design flexibility. The design policy for the BITF has been discussed in [18]. $\blacksquare$
Figure 1: Closed-loop system.
II-B BIBO stability of proper rational discrete-time IO SISO LTI system
In this study, we employ the standard definition of the BIBO stability [49]. Let $G\left(z\right)$ be a proper rational discrete-time IO SISO LTI transfer function. Then, the BIBO stability of $G\left(z\right)$ is defined as follows.
Definition 1.
Let $\left\{g_{k}\right\}_{k=0}^{\infty}$ be the impulse response sequence of $G\left(z\right)$ and let $\left\{u_{k}\right\}_{k=0}^{\infty}$ be an arbitrary bounded input sequence, i.e., there exists $0<\alpha_{u}<\infty$ such that $\left\lvert u_{k}\right\rvert<\alpha_{u}\;(k=0,1,2,\ldots)$. Then, $G\left(z\right)$ is said to be BIBO stable if there exists $0<\alpha_{y}<\infty$ satisfying $\left\lvert g_{k}\ast u_{k}\right\rvert<\alpha_{y}\;(k=0,1,2,\ldots)$. $\blacksquare$
The following facts are standard.
Proposition 1.
The transfer function $G\left(z\right)$ is BIBO stable if and only if one of the following conditions is satisfied:
  1. (C1)
    Its impulse response sequence $g=\left\{g_{k}\right\}_{k=0}^{\infty}$ is absolutely summable, i.e., $\left\lVert g\right\rVert_{1}<\infty$;
  2. (C2)
    All poles of $G\left(z\right)$ exist inside the unit circle. $\blacksquare$
(Proof.) See Section 4.3.2 in [49]. ∎
II-C Overview of the controller design problem in this study
Fig. 2 describes the control system considered in this study. In Fig. 2, the plant to be controlled has transfer function $P\left(z\right)$; $C\left(z;\theta\right)$ is the controller having the tunable parameter $\theta\in\Theta\subset\mathbb{R}^{n}$, where $\Theta$ is the search range for $\theta$. Here, $P\left(z\right)$ is a proper rational discrete-time IO SISO LTI system; $C\left(z;\theta\right)$ is a proper rational discrete-time IO SISO LTI system for any $\theta$. The class of the controller is a priori given by the designer. The setpoint reference signal, the control input, and the output of the actual system are denoted as $r$, $u$, and $y$, respectively. The reference signal $r$ is assumed to be bounded. The closed-loop transfer function from $r$ to $y$ is denoted as $T\left(z;\theta\right)=P\left(z\right)C\left(z;\theta\right)\left\{1+P\left(z\right)C\left(z;\theta\right)\right\}^{-1}$.
Here, the input data $u_{[0:N]}^{D}=\left\{u_{k}^{D}\right\}_{k=0}^{N}$ and the output data $y_{[0:N]}^{D}=\left\{y_{k}^{D}\right\}_{k=0}^{N}$ are collected by using the reference input $r_{[0:N]}^{D}=\left\{r_{k}^{D}\right\}_{k=0}^{N}\;\left(r_{0}^{D}\neq 0\right)$ and some initial controller $C\left(z;\theta_{0}\right)$ under the zero initial condition. In this study, we address the problem of finding $\theta$ such that $T\left(z;\theta\right)$ satisfies the designer-specified design specification including the iso-damping property, where $P\left(z\right)$ is unknown. Throughout this study, the input/output data is assumed to be noiseless. Under noisy conditions, the total variation denoising [50] is a practical choice to deal with the measurement noise.
Figure 2: Closed-loop system with unknown plant $P\left(z\right)$ and tunable controller $C\left(z;\theta\right)$.
III The proposed approach
This section describes the controller tuning problem described in Section II-C so as to achieve the iso-damping robust control.
III-A Robust controller design via MR-D3C framework
Fig. 3 sketches the proposed controller design approach based on the MR-D3C framework. Let $L_{flat}\left(s\right)$ be the continuous-time FO SISO LTI transfer function satisfying (2), and let $L_{flat}\left(s\right)$ be such that $L_{flat}\left(z\right)=\mathrm{c2d}\circ\mathrm{f2i}\left(L_{flat}\left(s\right)\right)$. In Fig. 3, $M_{ref}\left(z\right)=L_{flat}\left(z\right)\left\{1+L_{flat}\left(z\right)\right\}^{-1}$ is the FC-based reference model specifying the desired closed-loop characteristics. Note that $M_{ref}\left(z\right)$ has the iso-damping robustness owing to the specification of $L_{flat}\left(z\right)$. The designer a priori determines the specific choice of $L_{flat}\left(s\right)$. The meaning of $P\left(z\right)$, $C\left(z;\theta\right)$, $r$, $u$, and $y$ are the same as those in Fig 2. The following assumption can be easily satisfied by suitably designing $L_{flat}\left(z\right)$.
Assumption 1.
$L_{flat}\left(z\right)$ is proper and rational. $M_{flat}\left(z\right)$ has all poles inside the unit circle, implying $M_{flat}\left(z\right)$ is BIBO stable. $\blacksquare$
Then, the controller design problem in Section II-C can be formally formulated as follows:
Problem 1 (MR-D3C for iso-damping).
Find the optimal parameter $\theta^{\ast}\in\mathbb{R}^{n}$ of the controller on the basis of $u_{[0:N]}^{D}$ and $y_{[0:N]}^{D}$, such that $T\left(z;\theta^{\ast}\right)r_{[0:N]}^{D}$ is as close to $M_{ref}\left(z\right)r_{[0:N]}^{D}$ as possible in the mean square error sense. $\blacksquare$
Remark 3.
Since $L_{flat}\left(z\right)$ is constructed based on the FO transfer function, it is preferable that the class of the controller is also chosen as the FC-based controller such as $C\left(z;\theta\right)=\mathrm{c2d}\circ\mathrm{f2i}\left(C_{FOPID}\left(s;\theta\right)\right)$, where $C_{FOPID}\left(s;\theta\right)$ is the standard FO-PID controller with the tunable parameter $\theta$. This fact is demonstrated in Example 1 in Section IV. $\blacksquare$
Figure 3: MR-D3C approach to design the controller so as to achieve the iso-damping robustness.
III-B Fictitious reference signal
The proposed approach addresses Problem 1 based on the fictitious reference signal [37, 51]. The fictitious reference signal $\tilde{r}_{[0:N]}\left(\theta\right)=\left\{\tilde{r}_{k}\left(\theta\right)\right\}_{k=0}^{N}$ is defined as
$\tilde{r}_{[0:N]}\left(\theta\right)=\left(C\left(z;\theta\right)\right)^{-1}u_{[0:N]}^{D}+y_{[0:N]}^{D}.$ (4)
The fictitious reference signal has a unique property as follows:
Lemma 1.
For any $\theta$ such that $1+P\left(z\right)C\left(z;\theta\right)\neq 0$, $T\left(z;\theta\right)\tilde{r}_{[0:N]}\left(\theta\right)=y_{[0:N]}^{D}$. $\blacksquare$
(Proof) See Eq. (8) in [37]. ∎
Hereafter, we assume that $\tilde{r}_{0}\left(\theta\right)\neq 0$ for any $\theta$.
Remark 4.
One study has proposed a MR-D3C technique based on the fictitious reference signal: FRIT [37]. However, the controller provided by FRIT may destabilize the closed-loop system as pointed out in [50, 52]. To overcome this difficulty, several studies have examined to modify FRIT for IO control [50, 52]. $\blacksquare$
III-C The proposed controller design approach
In this study, we propose a novel approach for achieving iso-damping robust control in the MR-D3C framework. Specifically, to solve Problem 1, we propose a new MR-D3C technique: instability detecting FRIT for the iso-damping robust control (iso-IDFRIT).
Theorem 1 (iso-IDFRIT).
Let $M_{ref}\left(z\right)$ be the reference model constructed based on the FO open-loop transfer function with the flat-phase property, as stated in Section III-A. Let $m^{ref}=\left\{m_{k}^{ref}\right\}_{k=0}^{\infty}$ be the impulse response sequence of $M_{ref}\left(z\right)$ and $m_{[0:N]}^{ref}\coloneqq\left\{m_{k}^{ref}\right\}_{k=0}^{N}$. Then, the solution of Problem 1 is given by
[…]
$J\left(\theta\right)=\left(\left\lVert R^{D}\left(\tilde{R}\left(\theta\right)\right)^{-1}\mathrm{vec}\left(y_{[0:N]}^{D}\right)\right.\right.\$ (6)
where $R^{D}=\mathrm{Tpl}\circ\mathrm{vec}\left(r_{[0:N]}^{D}\right)$ and $\tilde{R}\left(\theta\right)=\mathrm{Tpl}\circ\mathrm{vec}\left(\tilde{r}_{[0:N]}\left(\theta\right)\right)$. $\blacksquare$ (Proof) Let $t\left(\theta\right)=\left\{t_{k}\left(\theta\right)\right\}_{k=0}^{\infty}$ be the impulse response sequence of $T\left(z;\theta\right)$. Due to Lemma 1,
$\mathrm{vec}\left(y^{D}\right)=\tilde{R}\left(\theta\right)\mathrm{vec}\left(t_{[0:N]}\left(\theta\right)\right)$ (7)
where $t_{[0:N]}\left(\theta\right)\coloneqq\left\{t_{k}\left(\theta\right)\right\}_{k=0}^{N}$. Since $\tilde{r}_{0}\left(\theta\right)\neq 0$, the impulse response of $T\left(z;\theta\right)$, which is the closed-loop transfer function from the reference input to the output when the controller parameter is $\theta$, from time 00 to $N$ is restored as
$\mathrm{vec}\left(t_{[0:N]}\left(\theta\right)\right)=\left(\tilde{R}\left(\theta\right)\right)^{-1}\mathrm{vec}\left(y_{[0:N]}^{D}\right).$ (8)
Therefore, $y_{[0:N]}\left(\theta\right)\coloneqq\left\{y_{k}\left(\theta\right)\right\}_{k=0}^{N}$, i.e., the output of $T\left(z;\theta\right)$ due to $r_{[0:N]}^{D}$, can be estimated as
$\mathrm{vec}\left(y_{[0:N]}\left(\theta\right)\right)$ $=R^{D}\mathrm{vec}\left(t_{[0:N]}\left(\theta\right)\right)$
$=R^{D}\left(\tilde{R}\left(\theta\right)\right)^{-1}\mathrm{vec}\left(y_{[0:N]}^{D}\right).$ (9)
On the other hand, the output of $M_{ref}\left(z\right)$ driven by $r_{[0:N]}^{D}$ can be computed as
$\mathrm{vec}\left(M^{ref}\left(z\right)r_{[0:N]}^{D}\right)=R^{D}\mathrm{vec}\left(m_{[0:N]}^{ref}\right)$ (10)
which completes the proof. ∎
Note that the minimization problem (5) is formulated in purely data-driven manner, i.e., the plant model $P\left(z\right)$ does not appear in (5). The desired closed-loop property specified by $M_{ref}\left(z\right)$, which has the iso-damping property, is realized by $C\left(z;\theta^{\ast}\right)$, since $\theta^{\ast}$ achieves the model matching between $M_{ref}\left(z\right)$ and $T\left(z;\theta^{\ast}\right)$ as closely as possible in the predetermined controller class. Therefore, the proposed approach, iso-IDFRIT, can design the desired controller $C\left(z;\theta^{\ast}\right)$ that achieves the iso-damping robustness without knowing the mathematical model $P\left(z\right)$ of the controlled plant.
Remark 5.
Equation (1) shows that the loss function $J\left(\theta\right)$ in the minimization problem (5) explicitly reflects the impulse response of $T\left(z;\theta\right)$. If the value of $J\left(\theta\right)$ is reasonably small, we can consider that $T\left(z;\theta\right)$ is BIBO stable. Contrarily, the controller making the value of $J\left(\theta\right)$ excessively large should destabilize the closed-loop system. That is, we can evaluate the BIBO stability of the resultant closed-loop system from a bounded reference input to a bounded output before implementing the controller. Therefore, iso-IDFRIT can design a reliable and practically feasible controller. $\blacksquare$
III-D Stability analysis
In this section, we analyze the BIBO stability of $T\left(z;\theta^{\ast}\right)$, i.e., the closed-loop system given by the proposed method, in an asymptotic case ($N\to\infty$). Here, we assume the following assumption:
Assumption 2.
The optimization problem in (5) is successfully solved for the asymptotic case. That is, $J\left(\theta\right)\to\alpha$ for $N\to\infty$, where $\alpha\in\mathbb{R}_{+}$ is a finite value. $\blacksquare$
The following lemma is the instrumental for the stability analysis:
Lemma 2.
Let $G\left(z\right)$ be a proper rational discrete-time IO SISO LTI transfer function. The impulse response sequence of $G\left(z\right)$ is denoted as $g=\left\{g_{k}\right\}_{k=0}^{\infty}$. Then, $G\left(z\right)$ is BIBO stable if and only if $\left\lVert g\right\rVert_{2}<\infty$. $\blacksquare$
(Proof) Let $(A,B,C,D)$ be the the minimal state-space realization of $G\left(z\right)$, i.e., $G(z)=C\left(zI-A\right)^{-1}B+D$ ($I$ is the identity matrix with the compatible dimension). Then, $g_{0}=D$ and $g_{k}=CA^{k-1}B$ for $k\geq 1$. Note that $\Lambda\left(A\right)$ coincides the poles of $G\left(z\right)$ including multiplicity due to the minimality of $(A,B,C,D)$ (Theorem 12.9.16 in [53]). This implies that $G\left(z\right)$ is BIBO stable if and only if $\rho\left(A\right)<1$ due to (C2) in Proposition 1.
Assume that $G\left(z\right)$ is BIBO stable. Then, due to (C1) in Proposition 1, $\left(\left\lVert g\right\rVert_{2}\right)^{2}<\left\lVert g\right\rVert_{\infty}\left\lVert g\right\rVert_{1}<\infty$. In contrast, suppose that $G\left(z\right)$ is not BIBO stable, i.e., $\rho\left(A\right)\geq 1$. Here, for $k\to\infty$, $\left\lvert g_{k}\right\rvert^{2}=\left\lvert CA^{k-1}B\right\rvert^{2}\to 0$ if and only if $A^{k-1}\to 0$. However, $A^{k-1}\to 0$ if and only if $\rho\left(A\right)<1$ (Theorem 5.6.12 in [54]), proving that $\left\lVert g\right\rVert_{2}$ is not finite. ∎
Here, the main result of the stability analysis is as follows:
Theorem 2.
Under the Assumptions 1 and 2, $T\left(z;\theta^{\ast}\right)$ is BIBO stable. $\blacksquare$
(Proof) Define
$\varepsilon_{[0:N]}\left(\theta^{\ast}\right)\coloneqq T\left(z;\theta^{\ast}\right)r_{[0:N]}^{D}-M_{ref}\left(z\right)r_{[0:N]}^{D}.$ (11)
Then,
. (12)
That is, Here, $J\left(\theta^{\ast}\right)=\left(\left\lVert\varepsilon_{[0:N]}\left(\theta^{\ast}\right)\right\rVert_{2}\right)^{2}=\left(\left\lVert\mathrm{vec}\left(\varepsilon_{[0:N]}\left(\theta^{\ast}\right)\right)\right\rVert_{2}\right)^{2}$ due to (8) and (LABEL:Eq_Vec-epsilon). Therefore,
$\begin{split}&\left\lVert\mathrm{vec}\left(t_{[0:N]}\left(\theta^{\ast}\right)\right)\right\rVert_{2}\\ &=\left\lVert\left(R^{D}\right)^{-1}\mathrm{vec}\left(\varepsilon_{[0:N]}\left(\theta^{\ast}\right)\right)+\mathrm{vec}\left(m_{[0:N]}^{ref}\right)\right\rVert_{2}\\ &\leq\beta_{RD}\sqrt{J\left(\theta^{\ast}\right)}+\left\lVert\mathrm{vec}\left(m_{[0:N]}^{ref}\right)\right\rVert_{2}\end{split}$ (14)
where $\beta_{RD}\in\mathbb{R}_{+}$ denotes the induced 2-norm of $\left(R^{D}\right)^{-1}$. Here, there exists a finite $\beta_{ref}\in\mathbb{R}_{+}$ such that $\lim_{N\to\infty}\left\lVert\mathrm{vec}\left(m_{[0:N]}^{ref}\right)\right\rVert_{2}=\left\lVert m^{ref}\right\rVert_{2}<\beta_{ref}$ due to Assumption 1 and Lemma 2. Finally,
$\begin{split}\lim_{N\to\infty}\left\lVert\mathrm{vec}\left(t_{[0:N]}\left(\theta^{\ast}\right)\right)\right\rVert_{2}&=\left\lVert t\left(\theta^{\ast}\right)\right\rVert_{2}\\ &\leq\sqrt{\alpha}\beta_{RD}+\beta_{ref}<\infty\end{split}$ (15)
which proves the conclusion. ∎
Remark 6.
In Theorem 2, Assumption 2 is a sufficient condition for proving the BIBO stability of $T\left(z;\theta^{\ast}\right)$; the BIBO stability may be achieved even if Assumption 2 does not hold. Therefore, the stability result in Theorem 2 suffers from some conservatism. Moreover, Theorem 2 is the asymptotic result. Analysis in the finite sample case and the derivation of the necessary and sufficient condition of the BIBO stability are left as future work. Nevertheless, the proposed loss function $J\left(\theta\right)$ in (6) explicitly reflects the pole information of $T\left(z;\theta^{\ast}\right)$ via its $\left(N+1\right)$-length impulse response $t_{[0:N]}\left(\theta\right)$. This feature allows us to infer that the resultant controller $C\left(z;\theta^{\ast}\right)$ stabilizes the closed-loop system from the reference input to the output in the BIBO sense, improving the reliability of the tuning result. $\blacksquare$
Remark 7.
We can interpret the loss function $J\left(\theta\right)$ from the viewpoint of the frequency domain. Specifically, since $J\left(\theta\right)=\sum_{n=1}^{N}\left\lvert\left(t_{k}\left(\theta\right)-m_{k}^{ref}\right)\ast r_{k}^{D}\right\rvert^{2}$,
where $\Phi_{rD}\left(e^{j\omega}\right)$ is the spectral density of the reference input $r_{[0:\infty]}^{D}$ used in the initial experiment. Therefore, the minimization of $J\left(\theta\right)$ corresponds to the minimization of the mismatch of the frequency responses $T\left(e^{j\omega};\theta\right)$ and $M_{ref}\left(e^{j\omega}\right)$ with the frequency weighting function $\Phi_{rD}\left(e^{j\omega}\right)$. $\blacksquare$
III-E Summary of the proposed controller design
The proposed data-driven design procedure can be summarized as follows. In Procedure 1, $C_{FO}\left(s;\theta\right)$ is an FO controller with the tunable parameter $\theta$. Procedure 1 can handle various types of controllers and reference models; the present approach is highly versatile and flexible.
  1. 1.
    Determine the control specifications including the iso-damping robustness. Determine the IO approximation and the discretization methods.
  2. 2.
    $L_{flat}\left(s\right)$ is given as the FO transfer function to to satisfy the specifications, and $L_{flat}\left(z\right)=\mathrm{c2d}\circ\mathrm{f2i}\left(L_{flat}\left(s\right)\right)$. The reference model is given as $M_{ref}\left(z\right)=L_{flat}\left(z\right)\left\{1+L_{flat}\left(z\right)\right\}$.
  3. 3.
    Define the class of controller $C\left(z;\theta\right)$ , e.g., $C\left(z;\theta\right)=\mathrm{c2d}\circ\mathrm{f2i}\left(C_{FO}\left(s;\theta\right)\right)$.
  4. 4.
    Set $C\left(z;\theta_{0}\right)$ , where $\theta_{0}$ is the initial guess of the controller parameter.
  5. 5.
    Acquire the initial data $u_{[0:N]}^{D}$ and $y_{[0:N]}^{D}$ through a closed-loop experiment using $C\left(z;\theta_{0}\right)$ and $r_{[0:N]}^{D}$.
  6. 6.
    Solve the optimization problem (5). The solution of (5) is denoted as $\theta^{\ast}$.
  7. 7.
    Return $C\left(z;\theta^{\ast}\right)$ as the resultant controller.
Procedure 1 Proposed approach (iso-IDFRIT).
Remark 8.
In practice, the IO approximation and discretization have to be carried out so as to implement an FO controller [17, 55]. The IO approximation and discretization may affect the closed-loop performance and stability [56]. Unlike conventional approaches, the proposed design scheme explicitly handles the IO-approximated and discretized controller. That is, the proposed approach evaluates the resultant closed-loop performance and stability for the IO-approximated and discretized, i.e., ready-to-implement, controller, showing the reliability of the proposed approach. $\blacksquare$
IV Numerical examples
[…]
IV-A Example 1: High-order process (FO control vs. IO control)
The controlled plant $P\left(z\right)$ is
$P\left(z\right)=\mathrm{c2d}\left(P_{3rd}\left(s\right)\right)$ (17)
where $P_{3rd}\left(s\right)$ is the following high-order process model [19]:
$P_{3rd}\left(s\right)=\frac{9}{\left(s+1\right)\left(s^{2}+2s+9\right)}.$ (18)
[…]
In this example, we demonstrate the superiority of FO control to IO control. Specifically, we compare the performance of the FO-PID controller $C_{FOPID}\left(z;\theta\right)=\mathrm{c2d}\circ\mathrm{f2i}\left(C_{FOPID}\left(s;\theta\right)\right)$ with that of the traditional IO-PID controller $C_{IOPID}\left(z;\theta\right)=\mathrm{c2d}\left(C_{IOPID}\left(s;\theta\right)\right)$. Here, $C_{IOPID}\left(s;\theta\right)$ and $C_{FOPID}\left(s;\theta\right)$ are as follows [60]:
$C_{IOPID}\left(s;\theta\right)=K_{p}+K_{i}\frac{1}{s}+K_{d}\frac{s}{1+\tau_{i}s}$ (19)
$C_{FOPID}\left(s;\theta\right)=K_{fp}+K_{fi}\frac{1}{s^{\lambda}}+K_{fd}\frac{s^{\mu}}{1+\tau_{f}s^{\mu}}$ (20)
where $K_{p},K_{i},K_{d},K_{fp},K_{fi},K_{fd},\lambda,\mu,\tau_{i},\tau_{f}\in\mathbb{R}_{+}\cup\left\{0\right\}$. The FOPID controller is a typical example of the FC-based control scheme [61]. In this study, we set $\tau_{i}=\tau_{f}=t_{s}$. The tuning conditions are listed in Table I. The IO approximation is performed via the Oustaloup recursive filter (order: 5; valid frequency range $\left(\omega_{b},\omega_{h}\right)=\left(10^{-4},10^{4}\right)\;\mathrm{rad/s}$) for both $C_{FOPID}\left(z;\theta\right)$ and $M_{ref}\left(z\right)$. In the initial data correction, the simulation time is $40\,\mathrm{s}$ and the reference input $r_{[0:N]}^{D}$ is the unit step. Note that the initial data for the controller tuning are the same to each other in the FO-PID controller and the IO-PID controller, since $C_{IOPID}\left(z;\theta_{0}\right)=C_{FOPID}\left(z;\theta_{0}\right)$. Fig. 4 shows the input/output data $u_{[0:N]}^{D}$ and $y_{[0:N]}^{D}$ for tuning both $C_{IOPID}\left(z;\theta\right)$ and $C_{FOPID}\left(z;\theta\right)$.
Table II summarizes the tuning results for $C_{IOPID}\left(z;\theta\right)$ and $C_{FOPID}\left(z;\theta\right)$ provided by the proposed approach (iso-IDFRIT). Table III lists the phase margin $\phi_{m}$ and the gain crossover frequency $\omega_{c}$ provided by the tuned controllers. Fig. 5 demonstrates the control results obtained by the resultant controllers. The Bode plots of $L_{flat}\left(z\right)$, $P\left(z\right)C_{IOPID}\left(z;\theta^{\ast}\right)$, and $P\left(z\right)C_{FOPID}\left(z;\theta^{\ast}\right)$ are shown in Fig. 6. In Fig. 5, the black line indicates the setpoint reference input. In Figs. 5 and 6, the green, blue, and red lines demonstrate the results due to the reference model ($M_{ref}\left(z\right)$ in Fig. 5 and $L_{flat}\left(z\right)$ in Fig. 6), $C_{IOPID}\left(z;\theta^{\ast}\right)$, and $C_{FOPID}\left(z;\theta^{\ast}\right)$, respectively. Tables II and III and Figs. 5 and 6 clearly show that the proposed approach successfully provides the IO- and FO-PID controllers yielding the desired closed-loop property. In particular, the closed-loop system due to the FO-PID controller achieves the closer specifications ($\phi_{m}$, $\omega_{c}$, the flat-phase) to the desired ones than that due to the IO-PID controller.
Fig. 7 demonstrates the robustness to the plant gain variation of the controllers tuned by the proposed approach. The black line indicates the setpoint reference input. The solid lines, the dotted lines, and the dashed lines represent the control results for $P\left(z\right)$ (i.e., the nominal plant), $1.5P\left(z\right)$ (i.e., the plant with increased gain), and $0.5P\left(z\right)$ (i.e., the plant with decreased gain), respectively. The gain crossover frequencies provided by $C_{IOPID}\left(z;\theta^{\ast}\right)$ are $1.3095\,\mathrm{rad/s}$ for $1.5P\left(z\right)$ and $5.3990\,\mathrm{rad/s}$ for $0.5P\left(z\right)$, whereas these due to $C_{FOPID}\left(z;\theta^{\ast}\right)$ are $1.4618\,\mathrm{rad/s}$ for $1.5P\left(z\right)$ and $5.2886\,\mathrm{rad/s}$ for $0.5P\left(z\right)$. Fig. 7 clearly shows that, although both $C_{IOPID}\left(z;\theta^{\ast}\right)$ and $C_{FOPID}\left(z;\theta^{\ast}\right)$ do not destabilize the closed-loop system, $C_{FOPID}\left(z;\theta^{\ast}\right)$ provides much stronger robustness of the overshoot amount to the plant gain variation, i.e., the iso-damping robustness, than $C_{IOPID}\left(z;\theta^{\ast}\right)$.
TABLE I: Conditions for Example 1
$\theta_{0}$ $\begin{bmatrix}K_{p}&K_{i}&K_{d}\\ \end{bmatrix}=\begin{bmatrix}1&0&0\\ \end{bmatrix}$
$\begin{bmatrix}K_{fp}&K_{fi}&\lambda&K_{fd}&\mu\\ \end{bmatrix}=\begin{bmatrix}1&0&1&0&1\\ \end{bmatrix}$
Search range $K_{p},K_{i},K_{d}\in\begin{bmatrix}0&5\\ \end{bmatrix}$
$K_{fp},K_{fi},K_{fd}\in\begin{bmatrix}0&5\\ \end{bmatrix}$,   $\lambda,\mu\in\begin{bmatrix}0&2\\ \end{bmatrix}$
$J\left(\theta_{0}\right)$ $9.8820\times 10^{2}$
TABLE II: Tuning results (Example 1)
$\theta^{\ast}$ IO $\begin{bmatrix}8.0397\times 10^{-1}&1.2125&3.3528\times 10^{-1}\\ \end{bmatrix}$
FO $\begin{bmatrix}1.3239&1.0370&1.1010&2.3253\times 10^{-1}&1.5465\\ \end{bmatrix}$
$J\left(\theta^{\ast}\right)$ IO $1.0172$
FO $4.2744\times 10^{-2}$
TABLE III: Frequency domain property provided by the tuned controllers (Example 1).
IO FO
Phase margin $\phi_{m}\,[\mathrm{deg.}]$ 73.3411 79.4165
Gain crossover frequency $\omega_{c}\,[\mathrm{rad/s}]$ 0.9371 1.0135
Figure 4: Initial input/output data for the proposed approach (Example 1)
Figure 5: Control results due to the tuned controllers (Example 1). The green and red lines almost overlap
Figure 6: Bode plots of $L_{flat}\left(z\right)$, $P\left(z\right)C_{IOPID}\left(z;\theta^{\ast}\right)$, $P\left(z\right)C_{FOPID}\left(z;\theta^{\ast}\right)$ (Example 1): (a) overview, and (b) enlarged phase plots around $1\,\mathrm{rad/s}$.
Figure 7: Robustness to the plant gain variation (Example 1).
IV-B Example 2: Soft robot (superiority over the conventional model-based approach)
One representative virtue of the proposed approach is the freedom from the necessity of the plant model and the model reduction. This feature allows the proposed approach to achieve better control performance and robustness than the reduced-model-based controller design scheme. Note that almost all conventional analytical and graphical design methodologies rely on the reduced order model, as discussed in Section I-B. We demonstrate the superiority and robustness of the proposed approach, compared to the traditional iso-damping controller design based on the reduced model. Throughout this example, the Oustaloup recursive filter is employed for the IO approximation (order: 7; valid frequency range $\left(\omega_{b},\omega_{h}\right)=\left(10^{-3},10^{6}\right)\;\mathrm{rad/s}$) for both controllers and reference models. We discretize the plant, the reference model, and the controller in the sampling time $t_{s}=10^{-2}\,s$.
The controlled plant $P\left(z\right)$ is
$P\left(z\right)=\mathrm{c2d}\left(P_{full}\left(s\right)\right)$ (21)
where $P_{full}\left(s\right)$ is as follows, which is the soft robot model (see [22]):
$P_{full}\left(s\right)=\frac{6}{s}\frac{54.893316s+2048.6337}{s^{2}+67.066887s+2048.7922}.$ (22)
[…]
$C_{FOPI}\left(s;\theta\right)=K_{fp}+K_{fi}\frac{1}{s^{\lambda}}$ (23)
where $K_{fp},K_{fi},\lambda\in\mathbb{R}_{+}\cup\left\{0\right\}$. The tuning condition is listed in Table IV. Fig. 8 describes the initial input/output data obtained using $C_{FOPI}\left(z;\theta_{0}\right)$, where the simulation time for initial data collection is $2\,\mathrm{s}$ and the reference input $r_{[0:N]}^{D}$ is the unit step. Note that the initial data is collected on the basis of the actual plant $P\left(z\right)$ in (21).
We compare the controller tuned by the proposed approach with that tuned by the iso-m method as the representative graphical tuning approach for FO controllers [22]. Specifically, based on the iso-m method [22], the controller $C_{FOPI}\left(s;\theta\right)$ is tuned using the following reduced plant model
$P_{reduced}\left(s\right)=\frac{6}{s}\frac{54.89}{s+54.89}$ (24)
[…]
We also compare the iso-damping robustness of $C_{iso-m}\left(z\right)$ and $C_{proposed}\left(z\right)$. Fig. 11 demonstrates the control results under the plant gain variation provided by $C_{iso-m}\left(z\right)$ and $C_{proposed}\left(z\right)$. Here, the black line indicates the setpoint reference input; the solid lines, the dotted lines, and the dashed lines represent the control results for $P\left(z\right)$ (i.e., the nominal plant), $1.5P\left(z\right)$ (i.e., the plant with increased gain), and $0.5P\left(z\right)$ (i.e., the plant with decreased gain), respectively. The gain crossover frequencies due to $C_{iso-m}\left(z\right)$ are $18.9349\,\mathrm{rad/s}$ for $1.5P\left(z\right)$ and $6.8977\,\mathrm{rad/s}$ for $0.5P\left(z\right)$; these provided by $C_{proposed}\left(z\right)$ are $17.0117\,\mathrm{rad/s}$ for $1.5P\left(z\right)$ and $6.9209\,\mathrm{rad/s}$ for $0.5P\left(z\right)$. As expected by the Bode plots in Fig. 10, the characteristic of the closed-loop system due to $C_{proposed}\left(z\right)$ is closer to the ideal iso-damping property than that due to $C_{iso-m}\left(z\right)$.
TABLE IV: Conditions for Example 2.
$\theta_{0}$ $\begin{bmatrix}K_{fp}&K_{fi}&\lambda\\ \end{bmatrix}=\begin{bmatrix}1&0&1\\ \end{bmatrix}$
Search range $K_{fp},K_{fi}\in\begin{bmatrix}0&15\\ \end{bmatrix}$,  $\lambda\in\begin{bmatrix}0&2\\ \end{bmatrix}$
$J\left(\theta_{0}\right)$ $3.8646\times 10^{1}$
TABLE V: Tuning result of the proposed approach (Example 2).
$\theta^{\ast}$ $\begin{bmatrix}8.8086\times 10^{-1}&3.8808&4.7498\times 10^{-1}\\ \end{bmatrix}$
$J\left(\theta^{\ast}\right)$ $2.0231\times 10^{-1}$
TABLE VI: Frequency domain property with the actual plant $P\left(z\right)$ (Example 2).
$C_{iso-m}$ $C_{proposed}$
Phase margin $\phi_{m}\,[\mathrm{deg.}]$ 67.7359 60.1370
Gain crossover frequency $\omega_{c}\,[\mathrm{rad/s}]$ 12.7008 12.0696
Figure 8: Initial input/output data for the proposed approach (Example 2).
Figure 9: Control results due to $C_{iso-m}\left(z\right)$ and $C_{proposed}\left(z\right)$ (Example 2). The green and red lines almost overlap
Figure 10: Bode plots of $L_{flat}\left(z\right)$, $P\left(z\right)C_{iso-m}\left(z;\theta^{\ast}\right)$, $P\left(z\right)C_{proposed}\left(z;\theta^{\ast}\right)$ (Example 2): (a) overview, and (b) enlarged phase plots around $12\,\mathrm{rad/s}$.
Figure 11: Robustness to the plant gain variation (Example 2).
V Discussion
In all examples, the proposed approach (iso-IDFRIT) successfully provides controllers realizing the desired closed-loop characteristics specified by the designer-defined reference model. The controller achieves the iso-damping robustness. In the examples, no closed-loop destabilization is observed, showing that the closed-loop system is BIBO stable from a bounded reference input to a bounded output. Therefore, we confirm the validity of the proposed approach.
Unlike traditional analytical and graphical approaches, the proposed approach does not require plant model experiments. Neither repeated experiments nor frequency response computations is necessary, whereas such procedures are indispensable for conventional data-driven and optimization-based approaches. In the present approach, the desired controller is automatically designed via solving the optimization problem defined based on only one-shot input/output data; the BIBO stability from a bounded reference input to a bounded output is explicitly considered for the IO-approximated and discretized (i.e., ready-to-implement) controllers. Hence, the proposed approach is a simple, practical, and reliable controller design technique so as to achieve the iso-damping property.
Example 1 shows not only the validity of the proposed approach but also the effectiveness of FO control for iso-damping robust control. Although both IO-PID and FO-PID controllers tuned by the present approach realize good control performance, the FO-PID controller outperforms the IO-PID controller from the viewpoint of the iso-damping robustness. This result is attributed to the fact that the reference model includes FO dynamics. Notably, the design burden of the FO-PID controller is not problematic compared with that of the IO-PID controller, owing to the simplicity of the proposed approach. The proposed approach contributes to the practical utilization of the virtues of FO control via reducing the implementation burden of FO controllers.
In Example 2, the FO-PI controller tuned by the present approach outperforms that tuned on the basis of the reduced order model. In other words, the proposed approach achieves better optimality than the model-based approach. This result stems from the gap between the actual plant and the model for controller design. In general, practical systems include high-order dynamics. Neglecting the high-order dynamics and designing the controller on the basis of the reduced-order model may adversely affect the control performance, robustness, and stability. On the other hand, the present approach can avoid such issues related to the unmodeled dynamics, since it relies on not the plant model but the experimentally collected data.
In this study, we assume that the reference model is given by the designer. That is, the designer must give the appropriate reference model. An important future direction may be the simplification of the design procedure for the reference model. For example, the flat-phase property is frequency-domain specification, whereas it is more straightforward and intuitive to design the reference model in the time-domain. The Auto-DDC approach [62] may be effective for addressing this issue. One study has proposed the data-driven tuning of parameters of both the controller and the reference model [63]. In the future, we will examine the Auto-DDC strategy and the simultaneous tuning technique of the controller and the reference model in iso-IDFRIT.
VI Conclusion
This study has presented novel controller design strategy to achieve iso-damping robust control. The proposed approach is simpler, more practical, and reliable than conventional iso-damping controller design methodologies. The proposed approach designs the iso-damping controller on the basis of the MR-D3C framework, requiring only one-shot input/output data. The plant model, the computation of the frequency response, or repeated experiments are unnecessary. The present controller tuning scheme explicitly evaluates the BIBO stability of the resultant closed-loop system from a reference input to an output. The numerical examples have demonstrated the effectiveness of the proposed approach. Moreover, this study has verified the advantage of FO control for iso-damping robust control.
In the future, we will analyze the stability and performance using the present approach in a finite sample under measurement noise.